# Fighting Exponentially Small Gaps by Counterdiabatic Driving

### v2

@@ A.2 Majorana fermions
One can rewrite
@@ A.3 Reflection parity symmetry
The Hamiltonian (1), and thus its Majorana representation (75), is odd under the chiral reflection parity symmetry $\widetilde{\Pi}_{\mathrm{R}}$ defined by Eq. (10), i.e. $\widetilde{\Pi}_{\mathrm{R}}\gamma_{k}\widetilde{\Pi}_{\mathrm{R}}=\gamma_{2L-k+1}$. Indeed, the couplings satisfy $J_{j}=J_{L-j}$, and one can write
@@
In the first line, the operator $\widetilde{\Pi}_{\mathrm{R}}$ was applied on the $\gamma$’s; in the second line, $\tilde{J}_{j}$ was exchanged with $\tilde{J}_{L-j}$; in the third line, it was set $l:=L-j$. $H$ is odd under the reflection parity $\widetilde{\Pi}_{\mathrm{R}}$, thus $\widetilde{\Pi}_{\mathrm{R}}$ is a chiral symmetry for $H$ [100].
@@ A.4 Reflection-symmetrized representation
Let us introduce symmetrized operators, as in Eqs. (12) in the main text. The inverse transformation is
@@
Let us split the Hamiltonian (75) into the different contributions: for $-\lambda J$ or $(1-\lambda)$, and bulk or boundary. It holds for $j=1,2,\dots,\ell$
@@
which is the same of Eq. (14) in the main text.
@@ A.5 Transformation of states
From the equations above, one can also find the relation between the vacuum state for the $c$’s and the one for the $\Gamma$’s. Defining |$\ket{0_{c}}$ to be the state annihilated by all $c_{j}$’s, and |$\ket{0_{\Gamma}}$ to be the one annihilated by all the $\Gamma_{k}^{-}$’s, it holds
|$\ket{0_{c}}=\prod_{k=1}^{\ell}\left(1+\Gamma_{2k-1}^{+}\Gamma_{2k}^{+}\right)\ket{0_{\Gamma}},$ (88)
@@ Appendix B Diagonalization of the effective hopping model
As a first thing, notice that for the choice of $J_{j}$’s in Eq. (2), the model is translation invariant in bulk (step 2), while only the sites $1$, $L-1$ and $L$ experience different couplings. This fact suggests that plane waves are eigenfunctions of the system, provided the boundary conditions are fixed.
@@ B.2 Left-localized state
When supplemented with Eqs. (93), (94) and the normalization condition, one obtains four equations for the four unknowns $a_{k}$,$b_{k}$,$k$,$\epsilon_{k}$. Here, we consider only the localized edge state, for which $k\equiv i\kappa$, $\kappa>0$. Eliminating $a_{k}/b_{k}$ and $\epsilon_{k}$ from Eq. (96) via Eqs. (93) and (94), one finds
@@ B.3 Right-localized state
$-B\psi_{L-2}+J_{\ell}\psi_{L}=\epsilon\psi_{L-1},\qquad J_{\ell}\psi_{L-1}-B\psi_{L}=\epsilon\psi_{L}.$ (99)
@@
$-Ba_{k}+Je^{ik}c_{k}=\epsilon_{k}b_{k},\qquad Jb_{k}e^{-ik}-Bc_{k}=\epsilon_{k}c_{k}.$ (101)
@@
Now, Eqs. (93), (94), (102) and the normalization condition are again four equations for the four unknowns $a_{k}$,$b_{k}$,$k$,$\epsilon_{k}$. The right-localized state will have an imaginary momentum $k\equiv-i\kappa$, $\kappa>0$, that can be determined by combining the cited equations. Since this would lead to cumbersome equations, it is convenient to utilize a workaround to determine the avoided crossing directly, as in Ref. [99].
@@ B.4 Avoided crossing
From Eq. (103) follows
@@
The last step consists in determining the magnitude of the gap at the avoided level crossing. Since $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ determined above are not eigenfunctions of the chain of finite length, one can argue that
@@
which corresponds to Eq. (19) in the main text. Indeed, the hybridization of $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ is approximately given by the off-diagonal matrix element $(\psi_{\mathrm{L}}|\mathcal{H}|\psi_{\mathrm{R}})$ in the reduced 2$\times$2 subspace spanned by $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ themselves. All the factors dropped in Eq. (108) are subleading with respect to the exponential term.
@@ C.1 Chiral reflection parity of the CD Hamiltonian
[…]
having set $y:=-x$. Thus, $H_{1}$ is even under the chiral reflection parity, Eq. (10), while $H$ is odd. The same conclusion can be reached by noticing that in Eq. (25), there is an even number of $H$’s at every order of the expansion.
@@
C.2 First order ansatz
@@
C.3 Second order ansatz
When considered together the first and second-order approximate CD Hamiltonians $H_{1}^{(1)}+H_{1}^{(2)}$, one can check that the variational equations, obtained from the minimization of the action Eq. (26), are
@@ C.3 Second order ansatz
[…]
One can check that the second order ansatz for the CD Hamiltonian, Eq. (30), is expressed in terms of Dirac fermions as
@@ E.1 Particle-number-nonconserving free fermions
$\mathcal{H}=\left(\begin{array}[]{c|c}\mathcal{H}^{+-}&\mathcal{H}^{++}\\ \hline\cr\mathcal{H}^{--}&\mathcal{H}^{-+}\end{array}\right).$ (146)
@@
$\mathcal{H}^{+-}=-\left(\mathcal{H}^{-+}\right)^{*},\qquad\mathcal{H}^{++}=-\left(\mathcal{H}^{--}\right)^{*},$ (147)
@@
$\mathcal{H}^{+-}=\left(\mathcal{H}^{+-}\right)^{\dagger},\qquad\mathcal{H}^{++}=-\left(\mathcal{H}^{++}\right)^{T}.$ (148)
@@
[…]
@@ E.2 Specification of the problem under consideration
One can check that the conjugation properties, Eq. (148), are satisfied. The first order CD Hamiltonian, Eq. (116), reads instead
@@
$\begin{cases}\displaystyle\Gamma^{+}_{k}=\frac{i\,\gamma_{k}+\gamma_{2L-k+1}}{\sqrt{2}}\,,&k=1,3,\dots,L\\ \displaystyle\Gamma^{+}_{k}=\frac{\gamma_{k}+i\,\gamma_{2L-k+1}}{\sqrt{2}}\,,&k=2,4,\dots,L-1\end{cases}\quad\implies\quad\begin{cases}\displaystyle\Gamma^{+}_{2j-1}=i\frac{c_{j}+c^{\dagger}_{j}+c_{L-j+1}-c^{\dagger}_{L-j+1}}{2},&j=1,\dots\ell+1\\ \displaystyle\Gamma^{+}_{2j}=i\,\frac{c_{j}-c^{\dagger}_{j}+c_{L-j+1}+c^{\dagger}_{L-j+1}}{2},&j=1,\dots\ell\end{cases}$ (165)
@@ E.3 Expectation values
$GS\displaystyle\expectationvalue{H[\lambda](t)}{\mathrm{GS}}$ $=-\lambda\sum_{j=1}^{L}J_{j}\expectationvalue{\sigma_{j}^{z}(t)\sigma_{j+1}^{z}(t)}{\rightarrow}-(1-\lambda)\sum_{j=1}^{L}\expectationvalue{\sigma_{j}^{x}(t)}{\rightarrow}$ (171)
=$0𝑐\displaystyle=-\lambda\sum_{j=1}^{L}\tilde{J}_{j}\expectationvalue{\left(c_{j}^{\dagger}(t)c_{j+1}^{\dagger}(t)+c_{j}^{\dagger}(t)c_{j+1}(t)-c_{j}(t)c_{j+1}^{\dagger}(t)-c_{j}(t)c_{j+1}(t)\right)}{0_{c}}$
@@
$0𝑐\displaystyle\sum_{j=1}^{L}\tilde{J}_{j}\expectationvalue{c_{j}^{\dagger}(t)c_{j+1}(t)}{0_{c}}$ $=\sum_{i,j=1}^{L}\tilde{J}_{j}\mathcal{V}_{j^{\dagger},i}\mathcal{V}_{j+1,i^{\dagger}},$ (174)
$0𝑐\displaystyle\sum_{j=1}^{L}\tilde{J}_{j}\expectationvalue{c_{j}(t)c_{j+1}^{\dagger}(t)}{0_{c}}$ $=\sum_{i,j=1}^{L}\tilde{J}_{j}\mathcal{V}_{j,i}\mathcal{V}_{(j+1)^{\dagger},i^{\dagger}},$ (175)
$0𝑐\displaystyle\sum_{j=1}^{L}\tilde{J}_{j}\expectationvalue{c_{j}(t)c_{j+1}(t)}{0_{c}}$ $=\sum_{i,j=1}^{L}\tilde{J}_{j}\mathcal{V}_{j,i}\mathcal{V}_{j+1,i^{\dagger}},$ (176)
$0𝑐\displaystyle\sum_{j=1}^{L}\expectationvalue{c_{j}^{\dagger}(t)c_{j}(t)}{0_{c}}$ $=\sum_{i,j=1}^{L}\mathcal{V}_{j^{\dagger},i}\mathcal{V}_{j,i^{\dagger}}.$ (177)
@@ Appendix F Derivation of the minimization equations for the Max Cut
where the ordering in the indices, $c_{1}<c_{2}<c^{\prime}_{2}$ of the connected interaction edges is fixed, and those denoted by $c^{\prime\prime}$ refer to the ones of the second order CD term. Here, the first term originates simply from the fact that in every three-body terms, the corresponding $\sigma^{x}_{j}$ will lead to either $\sigma^{z}_{j}$ or $\sigma^{y}_{j}$. The factor of three in the second line comes from the freedom of choosing one of the $c^{\prime\prime}_{i=1,2}$ spins in the construction. The third and fourth lines correspond to the cases when the $\sigma^{z}\sigma^{z}$ terms can potentially have non-zero commutator with the $\sigma^{y}\sigma^{x}$ and $\sigma^{x}\sigma^{y}$ terms.
The total second order term thus reads as
@@
+16($+16(1-\lambda)^{2}\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\beta^{2}_{c,c^{\prime}}+\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\left[4(1-\lambda)\beta_{c,c^{\prime}}-\lambda(\alpha_{c}+\alpha_{c^{\prime}})\right]^{2}+\lambda^{2}\sum_{c,c^{\prime},c^{\prime\prime}}\left(\delta_{c^{\prime}_{2},c_{1}}\beta^{2}_{c^{\prime},c}\delta_{c^{\prime}_{1},c^{\prime\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\beta^{2}_{c,c^{\prime}}\delta_{c^{\prime\prime}_{1},c^{\prime}_{2}}\right)$
@@
For computing the derivatives we again use combinatorial arguments. As all terms in $G\left(H^{(1)}_{1}+H^{(2)}_{1}\right)$ originating from different ordering of the connected spins are orthogonal to each other the constant prefactor in front of $\lambda^{2}\beta_{c,c^{\prime}}$ is given by four times the number of ways how a given spin can be involved in other interactions, i.e., by $12$. This can be seen easily by the fact, that one gets a non-zero operator for the commutator whenever the first two operators of $\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c^{\prime}_{2}}\sigma^{y}_{c^{\prime}_{3}}$ match any of the three in, $\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{3}$. As both of the resulting commutators give orthogonal operators, eventually one obtains two times the number times the given $c_{1},\,c_{2},\,c_{3}$ spins appear in the interactions. The off-diagonal, terms connecting different $\beta_{c,c^{\prime\prime}}$-s are obtained similarly as in the first order case, by enlisting all possible connections to other interaction pairs of $c,c^{\prime\prime}$ or $c^{\prime}c^{\prime\prime}$. One finds

### v3

@@ A.2 Majorana fermions
The Majorana representation is shown pictorially in Fig. 17(a). One can rewrite the Hamiltonian as
@@ A.3 Reflection parity symmetry
The Hamiltonian (1), and thus its Majorana representation (59), is odd under the chiral reflection parity symmetry $\widetilde{\Pi}_{\mathrm{R}}$, defined by
$\widetilde{\Pi}_{\mathrm{R}}\gamma_{k}\widetilde{\Pi}_{\mathrm{R}}=\gamma_{2L-k+1}.$ (60)
Indeed, the couplings satisfy $J_{j}=J_{L-j}$, and one can write
@@
In the first line, the operator $\widetilde{\Pi}_{\mathrm{R}}$ was applied on the $\gamma$’s; in the second line, $\tilde{J}_{j}$ was exchanged with $\tilde{J}_{L-j}$; in the third line, it was set l:=Ljl:=L-j. $H$ is odd under the reflection parity $\widetilde{\Pi}_{\mathrm{R}}$, thus $\widetilde{\Pi}_{\mathrm{R}}$ is a chiral symmetry for $H$ [127].
@@ A.4 Reflection-symmetrized representation
Let us introduce symmetrized operators
$\Gamma^{\pm}_{k}$ $=\frac{\pm i\gamma_{k}+\gamma_{2L-k+1}}{\sqrt{2}},\qquad$ $k=1,3,\dots,L,$ (67a)
$\Gamma^{\pm}_{k}$ $=\frac{\gamma_{k}\pm i\gamma_{2L-k+1}}{\sqrt{2}},\qquad$ $k=2,4,\dots,L-1,$ (67b)
where the factors $\pm i$ were chosen so that
$\{\Gamma_{k}^{+},\Gamma_{k^{\prime}}^{-}\}=\delta_{k,k^{\prime}},\qquad\{\Gamma_{k}^{+},\Gamma_{k^{\prime}}^{+}\}=\{\Gamma_{k}^{-},\Gamma_{k^{\prime}}^{-}\}=0,$ (68a)
$\widetilde{\Pi}_{\mathrm{R}}\Gamma_{k}^{\pm}\widetilde{\Pi}_{\mathrm{R}}=\pm i\Gamma_{k}^{\mp},\qquad(\Gamma^{\pm}_{k})^{\dagger}=\Gamma^{\mp}_{k}.$ (68b)
These correspond to Eqs. (7a) and (7b) in the main text, and are shown pictorially in Fig. 17(b). The inverse transformation is
@@
Let us split the Hamiltonian (59) into the different contributions: for $\lambda J$ or $(1-\lambda)$, and bulk or boundary. It holds for $j=1,2,\dots,\ell$
@@
which is the same of Eq. (6) in the main text, once the $J$’s are restored from the $\tilde{J}$’s. One can verify that $\widetilde{\Pi}_{\mathrm{R}}H[\lambda]\widetilde{\Pi}_{\mathrm{R}}=-H[\lambda]$ directly from Eqs. (7a)–(7b).
@@ A.5 Transformation of states
[…]
@@ Appendix B Diagonalization of the effective hopping model
As a first thing, notice that for the choice of $J_{j}$’s in Eq. (2), the model is translation invariant in bulk (step 2), while only the sites $1$, L1L-1 and $L$ experience different couplings. This fact suggests that plane waves are eigenfunctions of the system, provided the boundary conditions are fixed.
@@ B.2 Left-localized state
When supplemented with Eqs. (80), (81) and the normalization condition, one obtains four equations for the four unknowns $a_{k}$$b_{k}$$k\epsilon_{k}$. Here, we consider only the localized edge state, for which $k\equiv i\kappa$, $\kappa>0$. Eliminating $a_{k}/b_{k}$ and $\epsilon_{k}$ from Eq. (83) via Eqs. (80) and (81), one finds
@@ B.3 Right-localized state
$B\psi_{L-2}+J_{\ell}\psi_{L}=\epsilon\psi_{L-1},\qquad J_{\ell}\psi_{L-1}-B\psi_{L}=\epsilon\psi_{L}.$ (86)
@@
$Ba_{k}+Je^{ik}c_{k}=\epsilon_{k}b_{k},\qquad Jb_{k}e^{-ik}-Bc_{k}=\epsilon_{k}c_{k}.$ (88)
@@
Now, Eqs. (80), (81), (89) and the normalization condition are again four equations for the four unknowns $a_{k}$$b_{k}$$k\epsilon_{k}$. The right-localized state will have an imaginary momentum $k\equiv-i\kappa$, $\kappa>0$, that can be determined by combining the cited equations. Since this would lead to cumbersome equations, it is convenient to utilize a workaround to determine the avoided crossing directly, as in Ref. [99].
@@ B.4 Avoided crossing
The corresponding critical parameter value is given by $\lambda_{c}=1/(1+B_{c})$ changing in between $0.5$ and $0.6$ for the used values of $J^{\prime}$ and $J$ in the main text. From Eq. (90) follows
@@
The last step consists of determining the magnitude of the gap at the avoided level crossing. Since $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ determined above are not eigenfunctions of the chain of finite length, one can argue that
@@
which corresponds to Eq. (12) in the main text. Indeed, the hybridization of $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ is approximately given by the off-diagonal matrix element $(\psi_{\mathrm{L}}|\mathcal{H}|\psi_{\mathrm{R}})$ in the reduced $\times2$ subspace spanned by $|\psi_{\mathrm{L}})$ and $|\psi_{\mathrm{R}})$ themselves. All the factors dropped in Eq. (95) are subleading with respect to the exponential term.
@@ C.1 Chiral reflection parity of the CD Hamiltonian
$\widetilde{\Pi}_{\mathrm{R}}H_{1}\widetilde{\Pi}_{\mathrm{R}}$ $=-\frac{1}{2}\int_{-\infty}^{+\infty}dx\,f(x)\,\widetilde{\Pi}_{\mathrm{R}}e^{ixH}\partial_{\lambda}He^{-ixH}\widetilde{\Pi}_{\mathrm{R}}$ (96)
$=\frac{1}{2}\int_{-\infty}^{+\infty}dx\,f(x)\,e^{-ixH}\partial_{\lambda}He^{ixH}$ (97)
$=-\frac{1}{2}\int_{-\infty}^{+\infty}dy\,f(y)\,e^{iyH}\partial_{\lambda}He^{-iyH}=+H_{1},$ (98)
having set y:=xy:=-x. Thus, $H_{1}$ is even under the chiral reflection parity, Eq. (60), while $H$ is odd. The same conclusion can be reached by noticing that in Eq. (18), there is an even number of $H$’s at every order of the expansion.
@@
C.2 First-order ansatz
@@
The variational coefficients $\alpha_{j}$ are given by the solution of the system of equations obtained by the action minimization of Eq. (19),
$\left[8(1-\lambda)^{2}+\lambda^{2}\left(J_{j-1}^{2}+2J_{j}^{2}+J_{j+1}^{2}\right)\right]\alpha_{j}+2\lambda^{2}J_{j}\left(J_{j-1}\alpha_{j-1}+J_{j+1}\alpha_{j+1}\right)=-J_{j},$ (112)
with periodic boundary conditions implied, which can be solved via linear algebra techniques with a cost polynomial in $L$.
C.3 Second-order ansatz
When considered together, the first- and second-order approximate CD Hamiltonians $H_{1}^{(1)}+H_{1}^{(2)}$, one can check that the variational equations, obtained from the minimization of the action Eq. (19), are
@@ C.3 Second-order ansatz
$\alpha_{j}$ $\longleftrightarrow\,$ $\alpha_{L-j},\qquad$ $j=1,2,\dots,L$
$\beta_{j}$ $\longleftrightarrow$ $\beta_{L-j-1},\qquad$ $j=1,2,\dots,L-2$ (114)
$\beta_{L-1}$ $\longleftrightarrow$ $\beta_{L}.\qquad$
One can check that the second-order ansatz for the CD Hamiltonian, Eq. (22), is expressed in terms of Dirac fermions as
@@ E.1 Particle-number-nonconserving free fermions
[…]
@@
[…]
@@
[…]
@@
$O(t)=\widetilde{\mathrm{T}}e^{i\int_{0}^{t}H(t^{\prime})dt^{\prime}}O(0)\mathrm{T}e^{-i\int_{0}^{t}H(t^{\prime})dt^{\prime}},$ (143)
@@ E.2 Specification of the problem under consideration
One can check that the conjugation properties, Eq. (136), are satisfied. The first-order CD Hamiltonian, Eq. (103), reads instead
@@
$\begin{cases}\displaystyle\Gamma^{+}_{k}=\frac{i\,\gamma_{k}+\gamma_{2L-k+1}}{\sqrt{2}}\,,&k=1,3,\dots,L\\ \displaystyle\Gamma^{+}_{k}=\frac{\gamma_{k}+i\,\gamma_{2L-k+1}}{\sqrt{2}}\,,&k=2,4,\dots,L-1\end{cases}\quad\implies\quad\begin{cases}\displaystyle\Gamma^{+}_{2j-1}=i\frac{c_{j}+c^{\dagger}_{j}+c_{L-j+1}-c^{\dagger}_{L-j+1}}{2},&j=1,\dots\ell+1\\ \displaystyle\Gamma^{+}_{2j}=i\,\frac{c_{j}-c^{\dagger}_{j}+c_{L-j+1}+c^{\dagger}_{L-j+1}}{2},&j=1,\dots\ell.\end{cases}$ (153)
@@ E.3 Expectation values
where $H[\lambda=1](t)$ is the Hamiltonian corresponding to the parameter $\lambda$, evolved in the Heisenberg picture for a time $t$. Let us compute each term separately. Calling $\mathcal{V}:=\mathrm{T}e^{-2i\int_{0}^{t}\mathcal{H}(t^{\prime})dt^{\prime}}$, it holds
[…]
@@
[…]
@@ Appendix F Derivation of the minimization equations for the Max Cut
where the ordering in the indices, $c_{1}<c_{2}<c^{\prime}_{2}$ of the connected interaction edges is fixed, and those denoted by $c^{\prime\prime}$ refer to the ones of the second-order CD term. Here, the first term originates simply from the fact that in every three-body terms, the corresponding $\sigma^{x}_{j}$ will lead to either $\sigma^{z}_{j}$ or $\sigma^{y}_{j}$. The factor of three in the second line comes from the freedom of choosing one of the $c^{\prime\prime}_{i=1,2}$ spins in the construction. The third and fourth lines correspond to the cases when the $\sigma^{z}\sigma^{z}$ terms can potentially have non-zero commutator with the $\sigma^{y}\sigma^{x}$ and $\sigma^{x}\sigma^{y}$ terms.
The total second-order term thus reads as
@@
$+16(1-\lambda)^{2}\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\beta^{2}_{c,c^{\prime}}+\sum_{c,c^{\prime}}\delta_{c_{2},c^{\prime}_{1}}\left[4(1-\lambda)\beta_{c,c^{\prime}}-\lambda(\alpha_{c}+\alpha_{c^{\prime}})\right]^{2}+\lambda^{2}\sum_{c,c^{\prime},c^{\prime\prime}}\left(\delta_{c^{\prime}_{2},c_{1}}\beta^{2}_{c^{\prime},c}\delta_{c^{\prime}_{1},c^{\prime\prime}_{2}}+\delta_{c_{2},c^{\prime}_{1}}\beta^{2}_{c,c^{\prime}}\delta_{c^{\prime\prime}_{1},c^{\prime}_{2}}\right)$
@@
For computing the derivatives, we again use combinatorial arguments. As all terms in $G\left(H^{(1)}_{1}+H^{(2)}_{1}\right)$ originating from different ordering of the connected spins are orthogonal to each other the constant prefactor in front of $\lambda^{2}\beta_{c,c^{\prime}}$ is given by four times the number of ways how a given spin can be involved in other interactions, i.e., by $12$. This can be seen easily by the fact, that one gets a non-zero operator for the commutator whenever the first two operators of $\sigma^{z}_{c^{\prime}_{1}}\sigma^{x}_{c^{\prime}_{2}}\sigma^{y}_{c^{\prime}_{3}}$ match any of the three in $\sigma^{z}_{c_{1}}\sigma^{z}_{c_{2}}\sigma^{z}_{3}$. As both of the resulting commutators give orthogonal operators, eventually one obtains two times the number of times the given $c_{1},\,c_{2},\,c_{3}$ spins appear in the interactions. The off-diagonal terms connecting different $\beta_{c,c^{\prime\prime}}$-s are obtained similarly as in the first-order case, by enlisting all possible connections to other interaction pairs of $c,c^{\prime\prime}$ or $c^{\prime}c^{\prime\prime}$. One finds